Circular No. 7900

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SATELLITES OF JUPITER MPEC 2002-J54 provides 109 observations of 11 new outer satellites of

Jupiter. Following the communication to the Minor Planet Center by S. S. Sheppard and D. C. Jewitt, University of Hawaii, and J. Kleyna, University of Cambridge, of CCD observations with the 3.6-m Canada-France-Hawaii and 2.2-m University of Hawaii telescopes on pairs of nights in Dec. 2001 (Dec. 9–11 and 17–19), B. G. Marsden considered these 11 objects (in the mag range R = 22.2–23.5) likely to be satellites and provided ephemerides that allowed Sheppard’s measurement of further images on 2002 Jan. 8, as well as generally on another night during Jan. 9–16 and, with the help also of orbit calculations by R. Jacobson, Jet Propulsion Laboratory, during Feb. 10–11. In two cases Sheppard identified and measured images on Mar.

12. Marsden’s final perturbed orbit solutions (epoch 2002 May 6.0 TT), also given on MPEC 2002-J54, show all the satellites, designated S/2001 J 1–J 11, to have retrograde orbits (140^{o} < i < 166^{o} to the 2000.0 ecliptic) with P in the range 557–773 days and H from 15.4 to 16.8.

XTE J0929–314 D. K. Galloway, E. H. Morgan, R. A. Remillard and D. Chakrabarty,

Massachusetts Institute of Technology, report: “We have computed provisional orbital elements for XTE J0929–314 (cf. IAUC 7897), using pulse frequency measurements from RXTE/PCA observations between May 2.54 and 13.83 UT. Assuming a constant pulsar spin frequency and a circular orbit, our preliminary orbital solution has binary period 2614.75(15) s, projected semimajor axis 6.1(3) light-ms, and orbital epoch (time of 90^{o} mean longitude) 2002 May 11.4941(2) UT at the solar-system barycenter. The derived barycentric pulsar spin frequency is 185.1052(1) Hz. The inferred mass function of 2.7 × 10^{−7} M_{⊙} is the smallest measured for any stellar binary. For a 1.4-M_{⊙} neutron star, the minimum companion mass is 0.008 M_{⊙} (or 8.5 Jupiter masses). The x-ray source was active at a flux of 13 mCrab (2–10 keV) on May 15.69. RXTE observations are continuing.”

COMET C/2002 C1 (IKEYA-ZHANG) Visual m_{1} estimates: Apr. 28.85 UT, 4.4 (W. Hasubick, Buchloe, Ger-

many, 10×50 binoculars); May 1.92, 4.8 (R. J. Bouma, Groningen, The Netherlands, 7×50 binoculars); 6.98, 5.0 (A. Kammerer, Ettlingen, Germany, 9×63 binoculars); 11.24, 4.9 (A. Hale, Cloudcroft, NM, naked eye); 16.18, 5.1 (D. W. E. Green, Pack Monadnock Mtn., NH, naked eye).

2002 May 16  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green